# Linear Complexity Gibbs Sampling for Generalized Labeled Multi-Bernoulli Filtering

### v1

@@ 3       $n\sim\mathcal{U}(1,P)$;
Algorithm 4 $RGS^{+}$
III-D2 Deterministic-scan GS
The other extreme of complete randomness, namely deterministic coordinate selection can also be accommodated by $TGS^{+}$. The simplest deterministic scan is to traverse the coordinates according to the periodic sequence 1, 2, …, P, i.e., the selected coordinate at $t$-th iteration is given by $c(t)=1+\text{mod}(t\texttt{-}1,P)$. To realize this in the TGS framework, we consider a proposal of the form
$\phi_{i}(j|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))=\text{ }\delta_{c(t)}[i]\varphi_{i}(j|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))+\varepsilon(1-\delta_{c(t)}[i]),$
where $\varphi_{i}(j|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}})$ is given by (20), and $\varepsilon$ is a very small positive number. Note that the corresponding coordinate distribution is given by
$\rho(i|\gamma_{\text{+}})\propto\frac{\text{ }\delta_{c(t)}[i]\varphi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))}{\pi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))}+\frac{\varepsilon(1-\delta_{c(t)}[i])}{\pi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))},$
In principle, this special case has $\mathcal{O}(T(P+M))$ complexity. However, since the coordinate is effectively selected according to the prescribed sequence $c(t)$, the sampling step is not needed. Moreover, using the recursive construct in the initialization (Algorithm 3), it is possible to implement this so-called $DGS^{+}$ special case as described in Algorithm 5, which only incurs an $\mathcal{O}(TM)$ complexity. A similar approach based on the reverse scan order, i.e., $P$, $P-1$, …, 1, is possible by setting $c(t)=P-\text{mod}(t\texttt{-}1,P)$ in Algorithm 5 (line 5).
@@ III-D2 Deterministic-scan GS
Input : $\gamma_{\text{+}}^{(0)}$, $T$, $\alpha$, $\beta$, $[\eta_{i}(j)]_{i\text{=}1}^{P}$, $[\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))]_{i\text{=}1}^{P}$, $[\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$, $[\nu_{i}^{(\beta)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$
Output : $[\gamma_{\text{+}}^{(t)}]_{t\text{=}1}^{T}$
1   $P\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\text{=}1}^{P},\textsf{\footnotesize row})$;   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\text{=}1}^{P},\textsf{\footnotesize col})\texttt{-}2$;
2 foreach $t=1:T$ do
3     $n\leftarrow c(t)$;   $m\leftarrow c(t-1)$;
4     if $\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{m})>0$ then
5          $\widetilde{\pi}_{m}(\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{m})|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{m}}))\leftarrow 0$;
6
7     foreach $j=-1:M$ do
8          if $\widetilde{\pi}_{m}(j|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{m}}))=0$ $j\neq\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{n})$ then
9               $\widetilde{\pi}_{n}(j|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))\leftarrow 0$;
10
11          else
12               $\widetilde{\pi}_{n}(j|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))\leftarrow\eta_{n}(j)$;
13
14
15    $\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))\leftarrow\frac{\alpha\widetilde{\pi}_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))}{\nu_{n}^{(1)}\!(\gamma_{\text{+}}^{(t\text{-}1)})}+\frac{(1\text{-}\alpha)\widetilde{\pi}_{n}^{\beta}\!(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))}{\nu_{n}^{(\beta)}\!(\gamma_{\text{+}}^{(t\text{-}1)})}$;
16     $\gamma_{\text{+}}^{(t)}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1$:$M],\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}})))$;
17     $\gamma_{\text{+}}^{(t)}\leftarrow[\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{1:n\text{-}1}),\gamma_{\text{+}}^{(t)}\!(\ell_{n}),\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{n\text{+}1:P})]$;
18
Algorithm 5 $DGS^{+}$
Input : $\gamma_{\text{+}}^{(0)}$, $T$, $[\eta_{i}(j)]_{i\text{=}1}^{P}$, $[\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))]_{i\text{=}1}^{P}$, $[\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$
Output : $[\gamma_{\text{+}}^{(t)}]_{t\text{=}1}^{T}$
1   $P\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\text{=}1}^{P},\textsf{\footnotesize row})$;   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\text{=}1}^{P},\textsf{\footnotesize col})\texttt{-}2$;
2 foreach $t=1:T$ do
3     foreach $n=1:P$ do
4          $m\leftarrow 1+\textsf{\small mod}(P\texttt{-}2\texttt{+}n,P)$;
5          if $\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{m})>0$ then
6               $\widetilde{\pi}_{m}(\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{m})|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{m}}))\leftarrow 0$;
7
8          foreach $j=-1:M$ do
9               if $\widetilde{\pi}_{m}(j|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{m}}))=0$ $j\neq\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{n})$ then
10                    $\widetilde{\pi}_{n}(j|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))\leftarrow 0$;
11
12               else
13                    $\widetilde{\pi}_{n}(j|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))\leftarrow\eta_{n}(j)$;
14
15
16         $\gamma_{\text{+}}^{(t)}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1$:$M],\widetilde{\pi}_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}})))$;
17
18    $\gamma_{\text{+}}^{(t)}\leftarrow[\gamma_{\text{+}}^{(t)}\!(\ell_{1:P})]$;
19
Algorithm 6 $SGS^{+}$
@@ III-E TGS-GLMB Filter Implementation
  • SGS-based truncation is replaced by the proposed $\mathcal{O}(T(P+M))$ TGS algorithm (Algorithm 2);
@@
Thus, the resulting GLMB filter implementation incurs an overall complexity of $\mathcal{O}(T(P+M)+PM+T\log(T))$. The additional factor of $T\log(T)$ arises from the need to remove duplicates from the TGS output. Noting that in practice, the number $T$ of iterates is large, typically $T(P+M)>>PM$; nonetheless $\log(T)$ is small, and $P+M>>\log(T)$. This is a substantial reduction from the $\mathcal{O}(TP^{2}M)$ complexity of the SGS-based implementation.
IV Numerical Studies
Figure 3: Experiments for computation time.
This section presents numerical studies for GLMB filtering with truncation via the proposed TGS based schemes and the new recursive implementation of SGS.
  1. 1.
    $TGS^{+}$: $\mathcal{O}(T(P+M))$ TGS (Algorithm 2),
  2. 2.
    $RGS^{+}$: $\mathcal{O}(T(P+M))$ RGS (Algorithm 4),
  3. 3.
    $\overrightarrow{DGS}$${}^{+}$: $\mathcal{O}(TM)$ forward-scan DGS (Algorithm 5),
  4. 4.
    $\overleftarrow{DGS}$${}^{+}$: $\mathcal{O}(TM)$ backward-scan DGS (Algorithm 5),
  5. 5.
    $SGS^{+}$: $\mathcal{O}(TPM)$ SGS (Algorithm 6).
Note that, although all methods use the same initialization (Algorithm 3), only $TGS^{+}$ needs to compute $\rho(\cdot|\gamma_{\text{+}})$, whereas $RGS^{+}$/$SGS^{+}$ do not require the tempering related steps.
@@ IV Numerical Studies
A common linear Gaussian setup on a square $3$ $km$$\times$ $3$ $km$ surveillance region over 100 time steps is used throughout this section. The number of objects is time varying due to various births and deaths. The single-object state is a 4D vector $[x$ $\dot{x}$ $y$ $\dot{y}]^{\textsf{T}}$ of 2D position and velocity which follows a constant velocity model with sampling period $1$ $s$ and process noise standard deviation $\sigma_{p}=5$ $ms^{-2}$ on each axis. Existing objects have a constant survival probability of $P_{S}=0.99$. New objects appear according to an LMB birth density with $N_{B}=50$ components. Each LMB component has a fixed birth probability $r_{B,\texttt{+}}=0.01$ with a Gaussian birth density $p_{B,\texttt{+}}(\cdot)=\mathcal{N}(\cdot;m^{(i)},P_{B})$ for $i\in\{1,...,N_{B}\}$, where the means $m^{(i)}=[x^{(i)}$ 00 $y^{(i)}$ $0]^{\textsf{T}}$ are uniformly spaced in the $2.9$ $km$$\times$ $2.9$ $km$ square centered in the surveillance region, and the covariance is $P_{B}=\mathrm{diag}([10,10,10,10]^{\textsf{T}})^{2}$. The resulting mean rate of object births of $\lambda_{B}=0.5$ per scan, combined with the constant death or survival probability, results in a total of approximately $N_{X}=50$ trajectories, or a peak of approximately 33 objects simultaneously. Observations are noisy 2D positions with noise standard deviation $\sigma_{m}=10$ $m$ on each axis. All objects have a constant detection probability of $P_{D}=0.86$. Clutter follows a Poisson model with a mean rate of $\lambda_{c}=90$ returns per scan and a uniform density on the observation space. Each of the GS implementations are run for $T=5000$ iterations, and the $TGS^{+}$/$\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ use default mixture and tempering parameters of $\alpha=\beta=0.5$.
@@
The effects of individually varying key parameters affecting $T$, $P$, and $M$ are also examined. In addition to numerical studies with the default parameters described above, further experiments are also carried out according to Table II with varying number of GS iterations, total number of trajectories (or peak number of objects), plus detection and clutter rates.
Evaluations are undertaken for computation time, tracking accuracy, and (association) sample diversity, via (i) measured wall-clock times, (ii) OSPA [51] and OSPA${}^{(2)}$ [5] metrics, and (iii) the number of unique solutions per scan during GLMB filtering. Note that each Monte Carlo trial generates random births, motions, deaths (see the example in Fig. 4) and measurements according to the multi-object model parameters. All empirical averages are then reported over 350 Monte Carlo runs using MATLAB R2020a on a dual CPU machine with dual 64-Core AMD EPYC 7702 CPUs and 1TB RAM.
@@ IV-A Computation Time
Fig. 3 shows (in log scale) the measured wall-clock times averaged per scan (lower is better). It is imperative to examine both average and maximum times particularly in latency sensitive real-time applications [52]. As expected, the results in Figs. 3(a) and (b) show increasing computation times with increasing numbers of trajectories and iterations. The results in Figs. 3(c) and (d) show generally flat trends of computation times with varying rates of detection and clutter, since the plots use the same default values for the number of iterations and number of trajectories. In all cases, it can be seen that all linear complexity GS implementations have similar run times, but all are significantly faster than the higher complexity $SGS^{+}$. Of the linear complexity truncation strategies, $TGS^{+}$ is generally the most expensive, due to the additional calculations involved with tempering and coordinate selection. The deterministic scan $\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ have more similar run times to $RGS^{+}$, depending on the fraction of time spent in sampling versus overheads. Nonetheless, $TGS^{+}$/$\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ are observed to have improved tracking accuracy and improved sample diversity, when compared to $RGS^{+}$, as will be seen and further discussed in the next subsections.
Figure 5: Experiments for tracking accuracy.

### v2

@@ 3     $n\leftarrow c(t)$;   $m\leftarrow c(t-1)$;
11          else
12               $\widetilde{\pi}_{n}(j|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))\leftarrow\eta_{n}(j)$;
13
14
15    $\phi_{n}(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))\leftarrow\frac{\alpha\widetilde{\pi}_{n}(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))}{\nu^{(1)}_{n}\!(\gamma^{(t\text{-}1)}_{\text{+}})}+\frac{(1\text{-}\alpha)\widetilde{\pi}^{\beta}_{n}\!(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))}{\nu^{(\beta)}_{n}\!(\gamma^{(t\text{-}1)}_{\text{+}})}$;
16     $\gamma^{(t)}_{\text{+}}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1$:$M],\phi_{n}(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}})))$;
17     $\gamma^{(t)}_{\text{+}}\leftarrow[\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{1:n\text{-}1}),\gamma^{(t)}_{\text{+}}\!(\ell_{n}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n\text{+}1:P})]$;
18
Algorithm 5 $DGS^{+}$
@@
Input :  $\gamma^{(0)}_{\text{+}}$, $T$, $[\eta_{i}(j)]^{P}_{i\text{=}1}$, $[\widetilde{\pi}_{i}(\cdot|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))]^{P}_{i\text{=}1}$
Output : $[\gamma^{(t)}_{\text{+}}]^{T}_{t\text{=}1}$
1   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]^{P}_{i\text{=}1},\textsf{\footnotesize col})\texttt{-}2$;
2 foreach $t=1:T$ do
3     $\gamma^{(t)}_{\text{+}}\!(\ell_{P})\leftarrow\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{P})$;
4     foreach $n=1:P$ do
5          $m\leftarrow 1+\textsf{\small mod}(P\texttt{-}2\texttt{+}n,P)$;
6          if $\gamma^{(t)}_{\text{+}}\!(\ell_{m})>0$
7               $\widetilde{\pi}_{m}(\gamma^{(t)}_{\text{+}}\!(\ell_{m})|\gamma^{(t)}_{\text{+}}\!(\ell_{1\colon n\text{-}1}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n\text{+}1\colon P}))\leftarrow 0$;
8
9          foreach $j=-1:M$ do
10               if $\widetilde{\pi}_{m}(j|\gamma^{(t)}_{\text{+}}\!(\ell_{1\colon m\text{-}1}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{m\text{+}1\colon P}))=0$
11                                                       $j\neq\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n})$ $\widetilde{\pi}_{n}(j|\gamma^{(t)}_{\text{+}}\!(\ell_{1\colon n\text{-}1}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n\text{+}1\colon P}))\leftarrow 0$;
12
13               else
14                    $\widetilde{\pi}_{n}(j|\gamma^{(t)}_{\text{+}}\!(\ell_{1\colon n\text{-}1}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n\text{+}1\colon P}))\leftarrow\eta_{n}(j)$;
15
16
17         $\gamma^{(t)}_{\text{+}}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1$:$M],$                         $\widetilde{\pi}_{n}(\cdot|\gamma^{(t)}_{\text{+}}\!(\ell_{1\colon n\text{-}1}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n\text{+}1\colon P}))$;
18
19    $\gamma^{(t)}_{\text{+}}\leftarrow[\gamma^{(t)}_{\text{+}}\!(\ell_{1:P})]$;
20
Algorithm 6 $SGS^{+}$
@@ III-E TGS-GLMB Filter Implementation
  • SGS-based truncation is replaced by the proposed $\mathcal{O}(T(P+M))$ TGS algorithm (Algorithm 2); and
@@
Thus, the resulting GLMB filter implementation incurs an overall complexity of $\mathcal{O}(T(P+M)+PM+T\log T)$. The additional factor of $T\log T$ arises from the need to remove duplicates from the TGS output. Noting that in practice, the number $T$ of iterates is large, typically $T(P+M)>>PM$, nonetheless $\log T$ is small, and $P+M>>\log T$. This is a substantial reduction from the $\mathcal{O}(TP^{2}M)$ complexity of the SGS-based implementation.
IV Experimental Results
@@
This section presents numerical studies for GLMB filtering with truncation via the proposed TGS based schemes and the new recursive implementation of SGS:
  1. 1)
    $TGS^{+}$: $\mathcal{O}(T(P+M))$ TGS (Algorithm 2);
  2. 2)
    $RGS^{+}$: $\mathcal{O}(T(P+M))$ RGS (Algorithm 4);
  3. 3)
    $\overrightarrow{DGS}$${}^{+}$: $\mathcal{O}(TM)$ forward-scan DGS (Algorithm 5);
  4. 4)
    $\overleftarrow{DGS}$${}^{+}$: $\mathcal{O}(TM)$ backward-scan DGS (Algorithm 5);
  5. 5)
    $SGS^{+}$: $\mathcal{O}(TPM)$ SGS (Algorithm 6);
  6. 6)
    $RGS$: $\mathcal{O}(TPM)$ RGS (Subsection III-A); and
  7. 7)
    $SGS$: $\mathcal{O}(TP^{2}M)$ SGS (Algorithm 2a in [30]).
Note that, although all methods use the same initialization (Algorithm 3), only $TGS^{+}$ requires $\rho(\cdot|\gamma_{\text{+}})$, whereas $RGS^{+}/SGS^{+}$ do not require the tempering related steps. No parallel implementations are used in our experiments.
A common linear Gaussian setup on a square $3~{}km~{}\times~{}3~{}km$ surveillance region over 100 time steps is used throughout this section. The number of objects is time varying due to various births and deaths. The single-object state is a 4D vector $[x,\dot{x},y,\dot{y}]^{\textsf{T}}$ of 2D position and velocity which follows a constant velocity model with sampling period $1$ $s$ and process noise standard deviation $\sigma_{p}=5$ $ms^{-2}$ on each axis. Existing objects have a constant survival probability of $P_{S}=0.99$. New objects appear according to an LMB birth density with $N_{B}=50$ components. Each LMB component has a fixed birth probability $P_{B,\texttt{+}}(\ell_{i})=0.01$ and Gaussian birth density $f_{B,\texttt{+}}(x,\ell_{i})=\mathcal{N}(\cdot;m_{B}^{(i)},R_{B})$ for $i\in\{1,...,N_{B}\}$, where the means $m_{B}^{(i)}=[x^{(i)},0,y^{(i)},0]^{\textsf{T}}$ are uniformly spaced in the surveillance region, and $R_{B}=\mathrm{diag}([10,10,10,10]^{\textsf{T}})^{2}$. This implies a mean birth rate of $\lambda_{B}=0.5$ per scan, combined with the constant death or survival probability, results in a total of approximately $N_{X}=50$ trajectories, or a peak of approximately $33$ objects simultaneously. Observations are noisy 2D positions with noise standard deviation $\sigma_{m}=10$ $m$ on each axis. All objects have a constant detection probability of $P_{D}=0.86$. Clutter follows a Poisson model with a mean rate of $\lambda_{c}=90$ returns per scan and a uniform density on the observation space. Each of the GS implementations are run for $T=5000$ iterations, and the $TGS^{+}/\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ use default mixture and tempering parameters of $\alpha=\beta=0.5$.
The effects of individually varying key parameters affecting $T$, $P$, and $M$ are also examined. In addition to numerical studies with the default parameters described above, further experiments are also carried out according to Table II with varying number of GS iterations, total number of trajectories (or peak number of objects), plus detection and clutter rates. The effects of different mixture rates and tempering levels for $TGS^{+}$ are also studied.
@@
Mixture Rate $\alpha$ $0.1$ $0.9$ $(0.5)$
Tempering Level $\beta$ $0.1$ $0.9$ $(0.5)$
Evaluations are undertaken for computation time, tracking accuracy, and (association) sample diversity via (i) measured wall-clock times, (ii) OSPA [55] and OSPA${}^{(2)}$ [5] metrics, and (iii) the number of unique solutions per scan during GLMB filtering. Results on the tracking accuracy and sample diversity of $RGS/SGS$ are the same as per $RGS^{+}/SGS^{+}$ and hence omitted. Note that each Monte Carlo trial generates random births, motions, deaths (see the example in Fig. 3) and measurements according to the multi-object model parameters. All empirical averages are then reported over 1000 Monte Carlo runs using MATLAB R2020a on a dual CPU machine with dual 64-Core AMD EPYC 7702 CPUs and 1TB RAM.
Figure 4: Experimental results for computation time.
Figure 5: Experimental results for tracking accuracy.
@@ IV-A Computation Time
Fig. 4 shows (in log scale) the measured wall-clock times averaged per scan (lower is better). It is imperative to examine both average and maximum times particularly for latency sensitive real-time applications [56]. As expected, the results in Figs. 4(a) and (b) show increasing computation times with increasing numbers of trajectories and iterations. The results in Figs. 4(c) and (d) show generally flat trends of computation times with varying rates of detection and clutter, since the plots use the same default values for the number of iterations and number of trajectories. In all cases, it can be seen that all linear complexity GS implementations have similar run times, but all are significantly faster than the higher complexity methods $SGS^{+}/RGS/SGS$. Further, $RGS^{+}/SGS^{+}$ show better computational efficiency than $RGS/SGS$. Of the linear complexity truncation strategies, $TGS^{+}$ is generally the most expensive, due to the additional calculations involved with tempering and coordinate selection. The deterministic scan $\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ have more similar run times to $RGS^{+}$, depending on the fraction of time spent in sampling versus overheads. Nonetheless, $TGS^{+}/\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ are observed to have improved tracking accuracy and improved sample diversity, when compared to $RGS^{+}$, as will be seen and further discussed in the next subsections.